which is greater the anglr velocity of the hour hand of a watch or the angular velocity of the earth around its own axis?
1
Expert's answer
2012-09-18T12:13:01-0400
There are 365*4+1 = 1461 days in 4 years, so the Earth's angular velocity is

Ve = 1461/4 = 365.25 rounds per year.

The angular velocity of the hour hand of a watch is

Vh = 365 rounds per year.

So, the the angular velocity of the earth is greater than the angular velocity of the hour hand of a watch.